Consider the following realized annual returns:

Year

Stock A

Index

2000

23.6%

48.3%

2001

24.7%

28.7%

2002

30.5%

86.9%

2003

9.0%

23.1%

2004

-2.0%

0.2%

2005

-17.3%

-3.2%

2006

-24.3%

-27.0%

2007

32.2%

27.9%

2008

4.4%

-5.1%

2009

7.4%

-11.3%

a. Calculate the average of annual returns of the index.

b. Compute the standard deviation of annual returns of the index.

c. Compute the lower bound of the 95% confidence interval for annual returns of the index. Use the exact value from Excel, not an approximation.

d. Compute the geometric mean of annual returns of the index.

a. The average of annual returns of the index is %. (round to two decimals)

b. The standard deviation of annual returns of the index is %. (round to two decimals)

c. The lower bound of the 95% confidence interval of annual returns of the index is %. (round to one decimal)

d. The geometric mean of annual returns of the index is %. (round to two decimals)
Answer #1

a) Average of annual returns of the index = Sum of all returns / No of years=(48.3+28.7+89.9+23.1+0.2-3.2-27-27.9-5.1-11.3)/10= 168.50/10 = 16.85%.

b) Standard Deviation formula is as follows:

where,

= population standard deviation
= the size of the population
= each value from the population
= the population mean

Now let us tabulate some of the values for the formula

Index Index Mean Index – Mean Square of (Index- mean)
48.30% 16.85% 0.3145 989.1025
28.70% 16.85% 0.1185 140.4225
86.90% 16.85% 0.7005 4907.0025
23.10% 16.85% 0.0625 39.0625
0.20% 16.85% -0.1665 277.2225
-3.20% 16.85% -0.2005 402.0025
-27.00% 16.85% -0.4385 1922.8225
27.90% 16.85% 0.1105 122.1025
-5.10% 16.85% -0.2195 481.8025
-11.30% 16.85% -0.2815 792.4225

Mean of square of (Index-mean) = (989.1025+140.4225+4907.0025+39.0625+277.2225+402.0025+1922.8225+122.1025+481.8025+792.4225)/10 =1007.3965

Standard deviation = Square root of the mean of square of (Index-mean) = 31.73951 %. Rounding to two decimals standard deviation is  31.74%

Standard Error = Standard deviation / square root of number of observations = 31.74/ square root of 10 = 10.037

c) Lower bound at 95% confidence level = Historical Average Return - (2* Standard Error) = 16.85 - (2*10.037) = -3.224%

Lower bound at 95% confidence level rounded to one decimal = -3.2%

d) Geometric mean = Multiply all the annual returns and take the 10th root as there are annual returns for 10 years

=(48.30*28.7*86.9*23.1*0.2*(-3.2)*(-27)*27.9*(-5.1)*(-11.3)) ^(1/10)

=12.26946%

= 12.27% (rounded to two decimals)